OpenStudy (anonymous):

The radioactive isotope carbon-14 is present in small quantities in all life forms, and it is constantly replenished until the organism dies, after which it decays to stable carbon-12 at a rate proportional to the amount of carbon-14 present, with a half-life of 5556 years. Suppose C(t) is the amount of carbon-14 present at time t. The exponential decay differential equation that models this scenario is C'=-kC . Solve the differential equation to answer the following questions. (a) Find the value of the constant k in the differential equation. There is also Part B, if anybody can get (a) 1

OpenStudy (anonymous):

decay y=e^(-kt) growth y=e^(+kt) you have decay so C=e^(-kt) fits. C'=-k*e^(-kt)=-k*C also fits. now to find k: 1/2=e^(-k*5545) ln(1/2)=-k*5545 -ln(2)=-k*5545 k=ln(2)/5545 0.91=e^(-ln(2)*t/5545) ln(0.91)=-ln(2)*t/5545 t=(-ln(0.91)*5545/ln(2)

OpenStudy (anonymous):

was this ok for you? Reply if it was, :)

OpenStudy (anonymous):

sorry one sec

OpenStudy (anonymous):

Yes, Thank You!